OpenStudy (anonymous):

Suppose y varies inversely with x. Write an equation for the inverse variation.

OpenStudy (anonymous):

29. y= -9 when x=2

OpenStudy (anonymous):

y = k/x -9 = k/2 k = -18 y = -18/x x = -18/y